# Capturing Stable HDR Videos Using a Dual-Camera System

### v1

@@ V-B Experiments on self-captured videos
pre align fusion PSNR-$\mu$ PSNR-L SSIM-$\mu$ SSIM-L
GLA EFSM ACA $f^{guide}$
$\surd$ $\surd$ 41.29 39.47 0.9728 0.9888
$\surd$ $\surd$ $\surd$ 41.43 39.73 0.9722 0.9890
$\surd$ $\surd$ $\surd$ 41.17 39.33 0.9726 0.9780
$\surd$ $\surd$ $\surd$ $\surd$ 41.80 40.14 0.9731 0.9895
$\surd$ $\surd$ CA $\surd$ 41.61 39.80 0.9726 0.9892
$\surd$ $\surd$ $\surd$ 41.48 39.79 0.9723 0.9893
$\surd$ $\surd$ 41.13 39.12 0.9708 0.9887
40.81 38.70 0.9703 0.9784
V-C Ablation study
In this section, we investigate different components of our system individually to understand the capability of the system and the source of efficiency. Specifically, all quantitative evaluations are performed on Prabhakar’s dataset[30].
@@ V-C Ablation study
GLA PSNR-$\mu$ PSNR-$l$ SSIM-$\mu$ SSIM-$l$
9 $\surd$ $\surd$ $\surd$ 41.80 40.14 0.9731 0.9895
6 $\surd$ $\surd$ 41.17 39.33 0.9726 0.9780
6 $\surd$ $\surd$ 41.08 38.97 0.9729 0.9884
6 $\surd$ $\surd$ 41.19 39.54 0.9729 0.9882
@@
Proportions PSNR-$\mu$ PSNR-$l$ SSIM-$\mu$ SSIM-$l$
1:2 41.41 39.66 0.97281 0.97907
1:1 41.43 39.70 0.97284 0.98918
2:1 41.80 40.14 0.97311 0.98953
4:1 41.68 40.00 0.97251 0.98934
8:1 41.72 40.02 0.97255 0.98938
Asymmetric Cross-feature Fusion Subnetwork. As shown in Table V, replacing the asymmetric cross-feature fusion subnetwork with a generally used attention subnetwork [36] (line 7) results in a 0.67 dB drop in PSNR-$\mu$. Fig. 16 and Fig. 15 further confirms that our method effectively mitigates misalignment artifacts. Additionally, we evaluate the role of the asymmetric cross-attention (ACA) by comparing it with cross-attention (CA) mechanism. The ACA improves PSNR-$\mu$ by 0.19 dB (line 5). The cross-scale guidance features $f^{guide}$ further enhance fusion accuracy, reducing artifacts and contributing to a 0.32 dB gain in PSNR-$\mu$.
@@
Restoration Subnetwork In this section, we conduct ablation experiments to evaluate the role of the frequency-domain design and multiscale structure in HDR reconstruction. As shown in Table VIII, the baseline model (merging network in AHDNet) achieves a PSNR-$\mu$ of 40.81 and an SSIM-$\mu$ of 0.9719. First, we remove the frequency-domain design, and the results show that the PSNR-$\mu$ dropped by 0.09. Next, we perform an ablation experiment on the multiscale structure by removing it. The results indicate that, without the multiscale structure, the PSNR-$\mu$ decreased by 0.22. Our full model, which incorporates both the frequency-domain design and multiscale structure, yields the highest performance, achieving a PSNR-$\mu$ of 41.80 and an SSIM-$\mu$ of 0.97311.
TABLE VIII: Visual comparison of the separate effects of restoration network.
model PSNR-µ PSNR-L SSIM-µ SSIM-L
baseline 40.81 38.60 0.97192 0.98813
w/o mutil-scale 41.58 39.86 0.97249 0.98911
w/o DWT 41.71 40.04 0.97286 0.98937
Ours 41.80 40.14 0.97311 0.98953
@@ VI CONCLUSION
In this work, we propose a dual-camera system (DCS) to mitigate the temporal inconsistency with AE methods, and existing MEF methods can be seamlessly extended to our system with temporally consistent output. In addition, we introduce an exposure adaptive fusion network with two novel subnetworks to provide a more stable and efficient solution for video HDR reconstruction. Experimental results demonstrate that our approach outperforms existing methods, offering stable performance and high-quality reconstruction in complex real-world video environments. Our dual-camera framework provides a promising direction for real-time HDR video capture.

### v2

@@ V-D Ablation study
pre align fusion
Method GLA EFSM ACA $f^{guide}$ $\mu PSNR$-L $\mu SSIM$-L
1 41.29 39.47 0.9728 0.9888
2 41.43 39.73 0.9722 0.9890
3 41.17 39.33 0.9726 0.9780
4 CA 41.61 39.80 0.9726 0.9892
5 41.48 39.79 0.9723 0.9893
6 41.13 39.12 0.9708 0.9887
7 40.81 38.70 0.9703 0.9784
Ours 41.80 40.14 0.9731 0.9895
@@
GLA $\mu PSNR$-$l$ $\mu SSIM$-$l$
1 6 41.17 39.33 0.9726 0.9780
2 6 41.08 38.97 0.9729 0.9884
3 6 41.19 39.54 0.9729 0.9882
Ours 9 41.80 40.14 0.9731 0.9895
@@
Method Proportions $\mu PSNR$-$l$ $\mu SSIM$-$l$
1 1:2 41.41 39.66 0.9728 0.9791
2 1:1 41.43 39.70 0.9728 0.9892
3 4:1 41.68 40.00 0.9725 0.9893
4 8:1 41.72 40.02 0.9726 0.9894
Ours 2:1 41.80 40.14 0.9731 0.9895
Asymmetric Cross-feature Fusion Subnetwork. As shown in Table V, replacing the asymmetric cross-feature fusion subnetwork with a generally used attention subnetwork [56] (line 7) results in a 0.67 dB drop in PSNR-$\mu$. Fig. 14 and Fig. 13 further confirm that our method effectively mitigates misalignment artifacts. Additionally, we evaluate the role of the asymmetric cross-attention (ACA) by comparing it with cross-attention (CA) mechanism. The ACA improves PSNR-$\mu$ by 0.19 dB (line 5). The cross-scale guidance features $f^{guide}$ further enhance fusion accuracy, reducing artifacts and contributing to a 0.32 dB gain in PSNR-$\mu$.
@@
Restoration Subnetwork. In this section, we conduct ablation experiments to evaluate the role of the frequency-domain design and multi-scale structure in HDR reconstruction. As shown in Table VIII, the baseline model (merging network in AHDNet) achieves a PSNR-$\mu$ of 40.81 and an SSIM-$\mu$ of 0.9719. First, we remove the frequency-domain design, and the results show that the PSNR-$\mu$ dropped by 0.09. Next, we perform an ablation experiment on the multi-scale structure by removing it. The results indicate that, without the multi-scale structure, the PSNR-$\mu$ decreased by 0.22. Our full model, which incorporates both the frequency-domain design and multi-scale structure, yields the highest performance, achieving a PSNR-$\mu$ of 41.80 and an SSIM-$\mu$ of 0.97311.
@@
TABLE VIII: Comparison of the separate effects of the restoration subnetwork.
Method model $\mu PSNR$-L $\mu SSIM$-L
1 baseline 40.81 38.60 0.9719 0.9881
2 w/o multi-scale 41.58 39.86 0.9725 0.9891
3 w/o DWT 41.71 40.04 0.9729 0.9895
Ours full model 41.80 40.14 0.9731 0.9895
Figure 15: Performance variation of different models under simulated parallax. The horizontal axis denotes the introduced horizontal disparity in pixels, and the vertical axis shows the performance ($\mathrm{PSNR}$-$\mu$). All methods exhibit performance degradation as disparity increases, with SAFNet and DomainPlus showing the most significant drop. Our method consistently achieves the highest reconstruction quality across the entire disparity range, demonstrating superior robustness to parallax.
V-E limitation
Camera parallax. The inherent fixed-baseline nature of our dual-camera system inevitably introduces parallax between the two views, especially for objects at close range. Given our hardware configuration ($B=40\mathrm{mm}$, $f=6.5\mathrm{mm}$, $p=3.45\mu\mathrm{m}$, $f_{\mathrm{px}}\approx 1884,\mathrm{px}$), the expected horizontal disparity for a target at distance $Z$ follows
$\Delta x\approx\frac{f_{\mathrm{px}}B}{Z}.$ (13)
This corresponds to $\sim 9.4$ pixels for targets at $8\mathrm{m}$ and over $25$ pixels for those at $3\mathrm{m}$. Such disparity is inherent to our system geometry and can be mitigated by reducing the camera baseline.
We analyze the impact of parallax on model performance. As shown in Fig. 15, performance degrades as parallax increases. This trend is particularly pronounced for CNN-based approaches (SAFNet and DomainPlus), which are inherently sensitive to spatial misalignments due to their limited receptive fields and lack of explicit geometry modeling. Large disparities can cause feature mismatches across views, leading to ghosting, blur, or texture loss in the fused HDR output. Nevertheless, our method consistently maintains the highest reconstruction quality across all disparity ranges, demonstrating its robustness in the presence of parallax.
Efficiency Analysis The overall efficiency of HDR video processing depends on both data acquisition throughput and reconstruction latency. Our system addresses the acquisition bottleneck by employing an asynchronous dual-camera design, which decouples the capture of varying-exposure frames from the temporally consistent reference stream. This configuration eliminates the delays typically introduced by long high-exposure frames in single-camera systems, thereby enabling real-time HDR frame acquisition.
Although the proposed reconstruction model (EAFNet) is not explicitly optimized for real-time inference, it demonstrates competitive runtime performance in practice, as summarized in Table III. The system design thus provides a practical balance between capture efficiency and reconstruction quality, supporting high-rate video acquisition with high dynamic range and temporal consistency.
@@ VI CONCLUSION
In this work, we revisit the fundamental cause of temporal instability in alternating-exposure (AE) HDR video, which lies in the entanglement of temporal luminance anchoring with exposure-dependent detail selection, and propose a dual-stream paradigm that explicitly decouples these two roles. Specifically, we introduce an exposure-asymmetric dual-camera system (DCS) that fixes the reference exposure to stabilize frame-to-frame luminance while allowing the auxiliary stream to vary exposure for capturing high dynamic range details. To support this system, we design EAFNet to effectively suppress ghosting artifacts while preserving fine details in real-world scenes. Extensive experiments on multiple datasets and real-world sequences demonstrate that the proposed paradigm improves both temporal stability and reconstruction quality compared with AE-based baselines and HDR image deghosting methods, while remaining cost-efficient and deployment-friendly. Our dual-stream solution provides a promising direction for real-time HDR video capture.